17. At 30 °C Kw = 1.471 x 10 -14. What is the [ H] and [OH-] at a neutral pH? ( 12 pt)
Answer #1

We know for aqueous solution [ht] [ou] = kn - For neutral pH [++] = 104-1 [4+ ] [OH-] = kw = [+] [H ] = kw = 4+2 = kw => [++]